# Periodic nonlinear Schrödinger equation with distributional potential and invariant measures

### v1

@@ Abstract
In this paper, we continue some investigations on the periodic NLSE started by Lebowitz, Rose and Speer [16] and Bourgain [6] with the addition of a distributional multiplicative potential. We prove that the equation is globally wellposed for a set of data $\varphi$ of full normalized Gibbs measure, after suitable $L^{2}$-truncation in the focusing case. The set and the measure are invariant under the flow. The main ingredients used are Strichartz estimates on periodic NLS with distributional potential to obtain local well-posedness for low regularity initial data.
@@
$i\partial_{t}u=-\partial_{x}^{2}u+Vu+\lambda|u|^{p-2}u$ (1.1)
where $u$ is a function on $\mathbb{T}\times\mathbb{R}$ with $\mathbb{T}=\mathbb{R}/\mathbb{Z}$ for $p\geq 2$, $\lambda\in\mathbb{R}$ and $V$ a distributional potential. The difficulty of this equation lies in the roughness of the potential which makes the question of local well-posedness unclear. Indeed, the solution needs a priori to be regular enough for the product $Vu$ to make sense even for the linear equation
@@
with initial data $u(0)=u_{0}$. Our motivation comes from stochastic PDEs with $V=\xi$ the spatial white noise on $\mathbb{T}$, a random distribution in $C^{-\frac{1}{2}-\kappa}(\mathbb{T})$ for any $\kappa>0$. It corresponds to the derivative in the sense of distributions of a Brownian bridge and was first constructed by Paley and Zygmund [23, 24] as the random series
@@
where $(\xi_{k})_{k\geq 0}$ is a family of independent and identically distributed random variables of centered standard complex Gaussian and $\xi_{-k}=\overline{\xi_{k}}$. In this work, we consider a potential $V\in\mathcal{C}^{-1+\kappa}(\mathbb{T})$ for any $\kappa\in(0,1)$. This corresponds to the Young regime which does not involve any renormalization procedure and includes the particular case of white noise. Other examples of such distributionnal potential can be random potential with different correlation or highly oscillatory potential such as
@@
with a suitable growth condition on $(v_{n})_{n\geq 0}$ in order to ensure $V\in\mathcal{C}^{-1+\kappa}$. For the singular case, see for exemple [11] with $V$ a spatial white noise on $\mathbb{T}^{2}$ where the authors consider a regularization of the noise or [14, 18, 20] on the Anderson Hamiltonian using paracontrolled calculus. We follow the second approach which is a priori necessary in order to deal with low regularity initial data.
@@
is a priori doomed to fail for low regularity initial data due the singular product $Vu(s)$. For our purpose, we study the Hamiltonian
@@
which can be defined via its quadratic form $\mathbf{a}(u,v)=\langle\mathcal{H}u,v\rangle$ with form domain $H^{1}(\mathbb{T})$. One can prove that there exist an associated operator with domain
$\mathcal{D}(\mathcal{H})=\big{\{}u\in L^{2}(\mathbb{T})\ ;\ \exists v\in L^{2}(\mathbb{T}),\forall\varphi\in H^{1}(\mathbb{T}),a(u,\varphi)=\langle v,\varphi\rangle\big{\}}$ (1.7)
and discrete spectrum $(\lambda_{n})_{n\geq 0}$ with an orthonormal basis of eigenfunctions $(e_{n})_{n\geq 0}$. The paracontrolled calculus gives an explicit caracterization of the domain as
@@
with $X$ a solution to $-\partial_{x}^{2}X=V$ and $\mathsf{P}$ denotes Bony’s paraproduct constructed in [4] with the Paley-Littlewood decomposition, see Appendix A.
In order to study the nonlinear equation (1.1), one needs to understand the linear equation. The study of this equation without potential on a periodic domain goes back to Bourgain [6]. On the circle, the two major dispersive results on the free propagator are the inequalities
@@
for any $\varepsilon>0$ and $S_{N}$ the spectral projector on Fourier modes $|k|\leq N$. This has to be understood as a trade of integrability between time and space. Indeed, the solution $u(t)=e^{it\Delta}u_{0}$ of the linear equation with initial data $u_{0}\in L^{2}(\mathbb{T})$ has conserved mass hence $u\in L^{\infty}(\mathbb{R},L^{2}(\mathbb{T}))$ hence the trade of time integrability for space integrability, the second being at the cost of an arbitrary small positive Sobolev regularity for the initial data. This was used by Bourgain to prove local well-posedness for (1.1) for initial data in $H^{\sigma}$ with $\sigma\geq 0$ in the cubic case $p=4$ and $\sigma>0$ in the case $4<p\leq 6$. Following ideas from Burq, Gérard and Tzvetkov [7] and Mouzard and Zachhuber [20], we obtain Strichartz inequalities for the linear propagator associated to $\mathcal{H}$. This allows to use dispersive properties of the equation to get local well-posedness for low regularity initial data for equation (1.1). We believe that Strichartz inequalities with such general deterministic potential are of independent interest.
In order to extend local to global well-posedness, one usually relies on conserved quantity. For equation (1.1), the conserved quantities are the mass
@@
$\mathcal{E}(u)=\int_{\mathbb{T}}|\partial_{x}u(x)|^{2}\mathrm{d}x+\int_{\mathbb{T}}|u(x)|^{2}V(\mathrm{d}x)+\lambda\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x.$ (1.12)
The inital data needs to have finite energy in order to use these conservations, that is $u_{0}\in\mathcal{H}^{1}(\mathbb{T})$. For low regularity initial data, the only conserved quantity is the mass hence local well-posedness in $L^{2}(\mathbb{T})$ yields global solutions, this was proved for the cubic equation $p=4$ without potential by Bourgain [5] while in the case $4<p\leq 6$, he obtained local well-posedness in $H^{\sigma}(\mathbb{T})$ for $\sigma>0$. In the following work [6], he constructed global solution for random initial data distributed according to the Gibbs measure associated to the equation, using the invariance of the measure instead of conserved quantity. This is the path we follow here and prove global solution for random initial data given by the Gibbs measure. We insist on the fact that low regularity solution are challenging due to irregularity of the potential making it impossible to interpret equation (1.1) as a mild formulation associated to $-\partial_{x}^{2}$ even in the linear case.
@@
assuming for example that the operator is positive. Considering a basis of eigenfunctions with $(\lambda_{n})_{n\geq 1}$ and $(u_{n})_{n\geq 1}$ respectively the eigenvalues of $\mathcal{H}$ and the coefficients of a function $u$ in the basis, the measure can be rigorously interpreted at the product measure
$\mu(\mathrm{d}u)=\prod_{n\geq 1}\sqrt{\frac{\lambda_{n}}{\pi}}e^{-\lambda_{n}|u_{n}|^{2}}\mathrm{d}u_{n}$ (1.15)
assuming for example that the operator $\mathcal{H}$ is positive. We prove that this Gaussian measure associated to $\mathcal{H}$ is supported in $C^{\frac{1}{2}-\kappa}$ for any $\kappa>0$, using a comparaison with the case $V=0$ which correspond to the law of the Brownian bridge. In particular, the potential energy $\int_{\mathbb{T}^{d}}|u(x)|^{p}\mathrm{d}x$ makes sense for $\mu$-almost all functions $u$. In the defocusing case $\lambda>0$, the Gibbs measure
$\nu(\mathrm{d}u)=\frac{1}{Z}e^{-\lambda\int_{\mathbb{T}^{d}}|u(x)|^{p}\mathrm{d}x}\mu(\mathrm{d}u)$ (1.16)
is well-defined as the density is bounded by one. In the focusing case $\lambda<0$, this is not true anymore and one has to consider a cut-off. Since the norm $\|u(t)\|_{L^{2}(\mathbb{T})}$ is conserved, a natural family of measure is given by
$\nu_{B}(\mathrm{d}u)=\frac{\mathds{1}_{\|u\|_{L^{2}}\leq B}}{Z_{B}}e^{-\lambda\int_{\mathbb{T}^{d}}|u(x)|^{p}\mathrm{d}x}\mu(\mathrm{d}u)$ (1.17)
for any $B>0$. In the case $64\leq p<$ 6, we prove that this measure is invariant for generic cut-off parameter $B$ while a smallness assumption on $B$ is needed in the case $p=6$. This goes back to Lebowitz, Rose and Speer [16], see also [21] for a recent result for the critical parameter $B$ and Remark 4.3 for the relation with our measure.
We start with the description of the Hamiltonian $\mathcal{H}$ and its associated Schrödinger group $e^{it\mathcal{H}}$ using paracontrolled calculus in Section 2. In Section 3, we prove Strichartz inequalities which yields local well-posedness for low regularity initial data for equation (1.1). We also consider the truncated version of the equation with the spectral projector associated to $\mathcal{H}$ and prove the convergence of the solutions to the one of the untruncated equation. In Section 4, we construct the associated Gibbs measure, with a suitable cut-off in the focusing case, and prove global well-posedness on its support with invariance of the measure for $p\leq 6$. The Paley-Littlewood decomposition and the paraproduct are presented in Appendix A.
Acknoledgments : The first author benefits from the support of the French government “Investissements d’Avenir” program integrated to France 2030, bearing the following reference ANR-11-LABX-0020-01 and is partially funded by the ANR project ADA. The second author is grateful to Tristan Robert and Hugo Eulry for interesting discussions about this work.
@@
is well-defined with form domain $H^{1}(\mathbb{T})$, this is the content of the following proposition. In the following, we do not keep the space $\mathbb{T}$ in the notation since we always work on the circle.
@@
The derivative of a Brownian bridge $V=\mathrm{d}B\in C^{-\frac{1}{2}-\varepsilon}$ for any $\varepsilon>0$ corresponds to the Anderson Hamiltonian, this was its first construction by Fukushima and Nakao [13] on a finite segment. See also [17, 19] for a construction of the form in the singular case.
@@
see for example Ouhabaz’s book [22]. Moreover, the operator is self-adjoint, densely defined and bounded from below. Since $H^{1}$ is compactly imbedded in $L^{2}$, $\mathcal{H}$ has discrete spectrum $\lambda_{1}\leq\lambda_{2}\leq\ldots$ with an associated basis $(e_{n})_{n\geq 1}$ of $L^{2}$. Since the form domain of $\mathcal{H}$ and the Laplacian are the same, the first eigenvalue is simple, that is $\lambda_{1}<\lambda_{2}$ and there exists a positive ground state $\Psi\in\mathcal{D}(\mathcal{H})$, see [19, Theorem 4.1] for the details. The flow of the Schrödinger linear equation associated to $\mathcal{H}$ has the following spectral representation
@@
with initial data $u(0)=u_{0}$, the equation holding in the dual of the domain $\mathcal{D}(\mathcal{H})^{*}$. This is coherent with the case $V=0$ where the linear propagator solve the equation in $H^{-2}=(H^{2})^{*}$. However, the main difference here is that the domain $\mathcal{D}(\mathcal{H})$ does not contain smooth functions hence its dual $\mathcal{D}(\mathcal{H})^{*}$ is not a subspace of distributions and the equation is not satisfied in the sense of distributions. For the nonlinear equation (1.1), we will need a finer description of the operator as well as its domain.
@@
However, our application $\Gamma$ is not a unitary transformation of $L^{2}$ thus $\mathcal{H}^{\sharp}$ is not longer self-adjoint. The idea of such transformation was first used by Gubinelli, Ugurcan and Zachhuber [14] and Mouzard [18] to study the Anderson Hamiltonian in two dimensions. This was crucially used to obtain Strichartz inequalities by Mouzard and Zachhuber [20] as well as precise small time asymptotic and two-sided Gaussian bounds for the heat semigroup $e^{-t\mathcal{H}}$ by Bailleul, Dang and Mouzard [3]. We refer to Appendix A for the definitions of the tools from paracontrolled calculus such as the paraproduct $\mathsf{P}$ and the resonant product $\mathsf{\Pi}$.
@@
is continuous from $H^{\sigma}$ to itself for any $\sigma\leq 1+\kappa$ and it is invertible from $L^{2}$ to itself as a perturbation of the identity for $\|X\|_{C^{1+\kappa}}$ small enough. Using the truncated paraproduct $\mathsf{P}^{N}$ introduced in Appendix A, there exists $N(V)\geq 1$ such that $\|P_{u}^{N}X\|_{H^{1+\kappa}}\leq\frac{1}{2}\|u\|_{L^{2}}$ for $N\geq N(V)$ hence
@@
In particular, $\Gamma_{N}H^{1+\kappa}=H^{1+\kappa}\subset\mathcal{D}_{V}$ thus $\mathcal{D}_{V}$ is dense in $L^{2}$ and we have the explicit formula
@@
Proof : Let $u\in\mathcal{D}_{V}=\Gamma H^{2}$ and denote $v=\Phi(u)\in H^{2}$. The expression of $\mathcal{H}\Gamma v$ yields
@@
for a constant $c>0$ large enough. Using that $\Gamma$ is continuous from $L^{2}$ to itself, we get
$\|\Delta v+\mathcal{H}u\|_{L^{2}}\leq\frac{1}{2}\|v\|_{H^{2}}+c\|\Gamma\|_{L^{2}\to L^{2}}\|u\|_{L^{2}}$ (2.29)
which completes the proof.
@@
Instead of the paracontrolled calculus, one could use the exponential transform as in [15, 19] or the generalized Sturm-Liouville theory as in [12] and references therein to study the operator $\mathcal{H}$. These representation of the operator $\mathcal{H}$ implies a first order term which is problematic to obtain Strichartz inequalities. For example with the exponential transform, we have
@@
for $\sigma\in\mathbb{R}$. The previous results give the form domain $\mathcal{D}^{1}=H^{1}$ similar to the Laplacian while the domain $\mathcal{D}^{2}=\mathcal{D}_{V}$ depends on $V$.
@@
For any $|\sigma|\leq 1+\kappa$, we have $\mathcal{D}^{\sigma}=H^{\sigma}$.
@@
In order to prove Strichartz inequalities for $\mathcal{H}^{\sharp}$, the comparison with the Laplacian given by the following lemma is needed. This is exaclty where paracontrolled calculus comes into play and improve the naive bound
@@
allows to consider $u(t)=e^{-it\mathcal{H}}u_{0}$ as a solution in $L^{\infty}(\mathbb{R},L^{2})$ which satisfies (1.2) in $\mathcal{D}^{-2}$. In particular, the space $\mathcal{D}^{2}=\mathcal{D}(\mathcal{H})$ does not contain smooth functions thus $\mathcal{D}^{-2}=(\mathcal{D}^{2})^{*}$ is not a space of distributions. We want to use this approach to solve the nonlinear equation (1.1) for low regularity initial data with the mild formulation
$u(t)=e^{-it\mathcal{H}}u_{0}+\int_{0}^{t}e^{-i(t-s)\mathcal{H}}|u(s)|^{p-2}u(s)\mathrm{d}s$ (3.2)
with $u(0)=u_{0}\in H^{\sigma}$. Since $H^{\sigma}$ is an algebra for $\sigma>\frac{1}{2}$, this gives local well-posedness in $H^{\sigma}$ and global solution for $u\in H^{1}$ with the conservation of energy. In this section, we prove Strichartz inequalities for $\mathcal{H}$ to obtain local well-posedness for low regularity initial data, that is below the algebra condition. In the cubic case without potential, Bourgain obtained local well-posedness in $L^{2}$ thus global solution using the conservation of mass. Even in the cubic case, we have a positive critial threshold $\sigma_{c}>0$ which prevents us from obtaining global solution below $H^{1}$. Our argument is perturbative and relies on Bourgain’s result
@@
$u^{\sharp}(t)=e^{-it\mathcal{H}^{\sharp}}u_{0}+\int_{0}^{t}e^{-i(t-s)\mathcal{H}^{\sharp}}\Gamma^{-1}\big{(}|\Gamma u^{\sharp}(s)|^{p-2}\Gamma u^{\sharp}(s)\big{)}\mathrm{d}s$ (3.5)
@@
The loss of derivative here is similar in spirit to the one of Burq, Gérard and Tzvetkov [7] with the Laplacian on an arbitrary manifold and Mouzard and Zachhuber [20] with the Anderson Hamiltonian on a compact surfaces. This is due to the fact that our knowledge of the spectral properties of the operator is not precise enough to adapt Bourgain’s method based on Fourier’s series.
Proof : We work with the Paley-Littlewood projectors $\Delta_{j}$ on annulus of size $2^{j}$ in frequencies and consider $k,j\geq 0$, see Appendix A. For any $N\geq 1$, we have
@@
using the mild formulation for $e^{i(t-t_{n})\mathcal{H}^{\sharp}}$ with respect to $-\Delta$ hence

### v2

@@ Abstract
In this paper, we continue some investigations on the periodic NLSE started by Lebowitz, Rose and Speer [18] and Bourgain [6] with the addition of a distributional multiplicative potential. We prove that the equation is globally well-posed for a set of data $\varphi$ of full normalized Gibbs measure, after suitable $L^{2}$-truncation in the focusing case. The set and the measure are invariant under the flow. The main ingredients used are Strichartz estimates on periodic NLS with distributional potential to obtain local well-posedness for low regularity initial data.
MSC $\mathbf{2020}$ 35J10; 35Q55; 35A01; 35R60
Keywords – Schrödinger operator; Strichartz inequalities; Invariant measures; Paracontrolled calculus.
@@
$i\partial_{t}u=-\partial_{x}^{2}u+Vu+\lambda|u|^{m-2}u$ (1.1)
where $u$ is a function on $\mathbb{T}\times\mathbb{R}$ with $\mathbb{T}=\mathbb{R}/\mathbb{Z}$, $m\geq 2$ an integer, $\lambda\in\mathbb{R}$ and $V$ a distributional potential. The difficulty of this equation lies in the roughness of the potential which makes the question of local well-posedness unclear. Indeed, the solution needs a priori to be regular enough for the product $Vu$ to make sense even for the linear equation
@@
with initial data $u(0)=u_{0}$. Our motivation comes from stochastic PDEs with $V=\xi$ the spatial white noise on $\mathbb{T}$, a random distribution in the Besov-Hölder space $C^{-\frac{1}{2}-\kappa}(\mathbb{T})$ for any $\kappa>0$. It corresponds to the derivative in the sense of distributions of a Brownian bridge and was first constructed by Paley and Zygmund [25, 26] as the random series
@@
where $(\xi_{k})_{k\geq 0}$ is a family of independent and identically distributed random variables of centered standard complex Gaussian and $\xi_{-k}=\overline{\xi_{k}}$. In this work, we consider any potential $V\in\mathcal{C}^{-1+\kappa}(\mathbb{T})$ for any $\kappa\in(0,1)$. This corresponds to the Young regime which does not involve any renormalization procedure and includes the particular case of white noise. Other examples of such distributionnal potential can be random potential with different correlations or highly oscillatory potential such as
@@
with a suitable growth condition on $(v_{n})_{n\geq 0}$ in order to ensure $V\in\mathcal{C}^{-1+\kappa}(\mathbb{T})$. For the singular case beyond Young regime, see for exemple [12, 13, 29, 30] with $V$ a spatial white noise on $\mathbb{T}^{2}$ or $\mathbb{R}^{2}$ where the authors use an exponential transform or [16, 20, 22] which rely on a construction of the Anderson Hamiltonian using paracontrolled calculus. We follow the second approach to deal with low regularity initial data in a similar spirit to [22].
@@
is a priori doomed to fail for low regularity initial data due the singular product $Vu(s)$. This is in contrast with parabolic equation since the strong regularising properties of the heat semigroup are not available for the Schrödinger equation. For our purpose, we study the Hamiltonian
@@
which can be defined via its quadratic form $\mathbf{a}(u,v)=\langle\mathcal{H}u,v\rangle$ with form domain $H^{1}(\mathbb{T})$. One can prove that there exists an associated operator with domain
$\mathcal{D}(\mathcal{H})=\big{\{}u\in H^{1}(\mathbb{T})\ ;\ \exists v\in L^{2}(\mathbb{T}),\forall\varphi\in H^{1}(\mathbb{T}),a(u,\varphi)=\langle v,\varphi\rangle\big{\}}$ (1.7)
and discrete spectrum $(\lambda_{n})_{n\geq 1}$ with an orthonormal basis of eigenfunctions $(e_{n})_{n\geq 1}$. Paracontrolled calculus gives us an explicit caracterization of the domain as
@@
where $X$ is a solution to $-\partial_{x}^{2}X=V$ and $\mathsf{P}$ denotes Bony’s paraproduct constructed in [4] using Paley-Littlewood decomposition, see Appendix A. In particular, the spectral calculus associated to $\mathcal{H}$ yields solutions to associated linear equations such as (1.2).
A path to study the nonlinear equation (1.1) is to first understand the linear flow given by (1.2). The study of this equation without potential on a periodic domain goes back to Bourgain [6] motivated by Lebowitz, Rose and Speer [18]. On the circle, the two major dispersive results on the free propagator are the Strichartz inequalities
@@
for any $\varepsilon>0$ and $S_{N}$ the spectral projector on Fourier modes $|k|\leq N$. This has to be understood as a trade of integrability between time and space. Indeed, the solution $u(t)=e^{it\Delta}u_{0}$ of the linear equation with initial data $u_{0}\in L^{2}(\mathbb{T})$ has conserved mass hence $u\in L^{\infty}(\mathbb{R},L^{2}(\mathbb{T}))$ hence the trade of time integrability for spatial integrability, the second estimate being at the cost of an arbitrary small loss of positive Sobolev regularity for the initial data. This was used by Bourgain to prove local well-posedness for (1.1) for initial data in $H^{\sigma}$ with $\sigma\geq 0$ in the cubic case $m=4$ and $\sigma>0$ in the case $4<m\leq 6$. Following ideas from Burq, Gérard and Tzvetkov [7] and Mouzard and Zachhuber [22], we obtain Strichartz inequalities for the linear propagator associated to $\mathcal{H}$. This allows to use dispersive properties of the equation to get local well-posedness for low regularity initial data for equation (1.1). We believe that Strichartz inequalities with such general deterministic rough potential are of independent interest. For example, this allows to generalize without loss the second estimate obtained by Bourgain for a potential $V\in L^{\infty}(\mathbb{T})$.
In order to extend local to global well-posedness, one usually relies on conserved quantity. For equation (1.1), the two principal conserved quantities are the mass
@@
$\mathcal{E}(u)=\int_{\mathbb{T}}|\partial_{x}u(x)|^{2}\mathrm{d}x+\int_{\mathbb{T}}|u(x)|^{2}V(\mathrm{d}x)+\lambda\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x.$ (1.12)
The inital data needs to have finite energy in order to use its conservation, that is $u_{0}\in\mathcal{H}^{1}(\mathbb{T})$. For low regularity initial data, the only conserved quantity is the mass hence local well-posedness in $L^{2}(\mathbb{T})$ yields global solutions, this was proved for the cubic equation $m=4$ without potential by Bourgain [5] while in the case $4<m\leq 6$, he obtained local well-posedness in $H^{\sigma}(\mathbb{T})$ only for $\sigma>0$. In his following work [6], he constructed global solution for random initial data distributed according to the Gibbs measure associated to the equation using the invariance of the measure instead of conserved quantity. We obtain a similar result here and construct global solutions for random initial data given by the Gibbs measure. We insist on the fact that low regularity solutions are challenging due to irregularity of the potential making it impossible to interpret equation (1.1) as a mild formulation associated to $-\partial_{x}^{2}$ even in the linear case. Our method for the globalisation differs from Bourgain [6] and seems new in the context of dispersive PDEs. It was first used in the parabolic case by Da Prato and Debussche [10].
@@
assuming for example that the operator is positive. Considering a basis of eigenfunctions with $(\lambda_{n})_{n\geq 1}$ and $(u_{n})_{n\geq 1}$ respectively the eigenvalues of $\mathcal{H}$ and the coefficients of a function $u$ in the basis, the measure can be rigorously interpreted as the product measure
$\mu(\mathrm{d}u)=\prod_{n\geq 1}\sqrt{\frac{\lambda_{n}}{\pi}}e^{-\lambda_{n}|u_{n}|^{2}}\mathrm{d}u_{n}.$ (1.15)
We prove that this Gaussian measure associated to $\mathcal{H}$ is supported in $C^{\frac{1}{2}-\kappa}$ for any $\kappa>0$, which is natural since the case $V=0$ corresponds to the law of the Brownian bridge. In particular, the potential energy $\int_{\mathbb{T}^{d}}|u(x)|^{m}\mathrm{d}x$ makes sense for $\mu$-almost all functions $u$. In the defocusing case $\lambda>0$, the Gibbs measure
$\nu(\mathrm{d}u)=\frac{1}{Z}e^{-\lambda\int_{\mathbb{T}^{d}}|u(x)|^{m}\mathrm{d}x}\mu(\mathrm{d}u)$ (1.16)
is well-defined as the density is bounded by one and $Z<\infty$. In the focusing case $\lambda<0$, this does not hold anymore and one has to consider a cut-off. Since the norm $\|u(t)\|_{L^{2}(\mathbb{T})}$ is conserved, a natural family of measure is given by
$\nu_{B}(\mathrm{d}u)=\frac{\mathds{1}_{\|u\|_{L^{2}}\leq B}}{Z_{B}}e^{-\lambda\int_{\mathbb{T}^{d}}|u(x)|^{m}\mathrm{d}x}\mu(\mathrm{d}u)$ (1.17)
for any $B>0$. In the case $64\leq m<$ 6, we prove that this measure is invariant for generic cut-off parameter $B$ while a smallness assumption on $B$ is needed in the case $m=6$. The introduction of a mass cut-off goes back to Lebowitz, Rose and Speer [18], see also [23] for a recent result for the critical parameter $B$ in the case $m=6$ and Remark 4.3 for the relation with our measure.
In Section 2, we study the Hamiltonian $\mathcal{H}$ and give the first properties of its associated Schrödinger group $e^{it\mathcal{H}}$ using paracontrolled calculus. In Section 3, we prove Strichartz inequalities which yields local well-posedness for low regularity initial data for equation (1.1). We also consider the truncated version of the equation with the spectral projector associated to $\mathcal{H}$ and prove the convergence of the solutions to the untruncated equation. In Section 4, we construct the associated Gibbs measure with a suitable cut-off in the focusing case and prove global well-posedness on its support with invariance of the measure for $m\leq 6$. The Paley-Littlewood decomposition and the paraproduct are presented in Appendix A.
Acknoledgments : The first author benefits from the support of the French government “Investissements d’Avenir” program integrated to France 2030, bearing the following reference ANR-11-LABX-0020-01 and is partially funded by the ANR project ADA. The second author is grateful to Tristan Robert, Hugo Eulry and Nicolas Camps for interesting discussions about this work.
@@
is well-defined for any $u,v\in H^{1}(\mathbb{T})$, this is the content of the following proposition. In the following, we do not keep the space $\mathbb{T}$ in the notation since we always work on the circle.
@@
Considering the derivative of a Brownian bridge $V=\mathrm{d}B\in C^{-\frac{1}{2}-\varepsilon}$ for any $\varepsilon>0$ corresponds to the Anderson Hamiltonian, this was its first construction by Fukushima and Nakao [15] on a finite segment. See also [19, 21] for a construction of the form in the singular case.
@@
see for example Ouhabaz’s book [24]. Moreover, the operator is self-adjoint, densely defined and bounded from below. Since $H^{1}$ is compactly imbedded in $L^{2}$, $\mathcal{H}$ has discrete spectrum $\lambda_{1}\leq\lambda_{2}\leq\ldots$ with an associated basis $(e_{n})_{n\geq 1}$ of $L^{2}$. Note that since the form domain of $\mathcal{H}$ and of the Laplacian are the same, the first eigenvalue is simple, that is $\lambda_{1}<\lambda_{2}$ and there exists a positive ground state $\Psi\in\mathcal{D}(\mathcal{H})$, see [21, Theorem 4.1] for the details. The flow of the Schrödinger linear equation associated to $\mathcal{H}$ has the following spectral representation
@@
with initial data $u(0)=u_{0}$, the equation being interpreted in the dual of the domain $\mathcal{D}(\mathcal{H})^{*}$. This is coherent with the case $V=0$ where the linear propagation of any $u_{0}\in L^{2}$ solves the equation in $H^{-2}=(H^{2})^{*}$. However, the main difference here is that the domain $\mathcal{D}(\mathcal{H})$ does not contain smooth functions hence its dual $\mathcal{D}(\mathcal{H})^{*}$ is not a subspace of distributions and the equation is not satisfied in the sense of distributions. For the nonlinear equation (1.1), we will need a finer description of the operator as well as its domain.
@@
However the map $\Gamma$ is not a unitary transformation of $L^{2}$ thus $\mathcal{H}^{\sharp}$ is not longer self-adjoint. The idea of such transformation was first used by Gubinelli, Ugurcan and Zachhuber [16] and Mouzard [20] to study the Anderson Hamiltonian in two dimensions. This was crucially used to obtain Strichartz inequalities by Mouzard and Zachhuber [22] as well as precise small time asymptotic and two-sided Gaussian bounds for the heat semigroup $e^{-t\mathcal{H}}$ by Bailleul, Dang and Mouzard [3]. We refer to Appendix A for the definitions of the tools from paracontrolled calculus such as the paraproduct $\mathsf{P}$ and the resonant product $\mathsf{\Pi}$.
@@
is continuous from $H^{\sigma}$ to itself for any $\sigma\leq 1+\kappa$ and it is invertible from $L^{2}$ to itself as a perturbation of the identity for $\|X\|_{C^{1+\kappa}}$ small enough. Using the truncated paraproduct $\mathsf{P}^{N}$ introduced in Appendix A which corresponds to a truncation of the low frequencies, there exists $N(V)\geq 1$ such that $\|P_{u}^{N}X\|_{H^{1+\kappa}}\leq\frac{1}{2}\|u\|_{L^{2}}$ for $N\geq N(V)$ hence
@@
In particular, the continuity of $\Gamma_{N}$ gives that the domain is dense in $H^{1+\kappa}$ thus in $L^{2}$.
Corollary 2.4.
The space $\mathcal{D}_{V}$ is dense in $H^{1+\kappa}$.
Proof : Let $f\in H^{1+\kappa}$. Then
$f=(\Gamma_{N}\circ\Phi_{N})(f)=\Gamma_{N}g$ (2.22)
with $g=\Phi_{N}(f)\in H^{1+\kappa}$. Since $H^{2}$ is dense in $H^{1+\kappa}$, there exists a familly $(g_{\varepsilon})_{\varepsilon}\subset H^{2}$ such that
$\lim_{\varepsilon\to 0}\|g-g_{\varepsilon}\|_{H^{1+\kappa}}=0.$ (2.23)
Since $\Gamma_{N}$ is continuous from $H^{1+\kappa}$ to itself, we get
$\lim_{\varepsilon\to 0}\|f-\Gamma_{N}g_{\varepsilon}\|_{H^{1+\kappa}}=0.$ (2.24)
$\square$
We have the explicit formula
@@
Proof : Let $u\in\mathcal{D}_{V}=\Gamma H^{2}$ and denote $v=\Phi(u)\in H^{2}$. The expression of $\mathcal{H}u=\mathcal{H}\Gamma v$ yields
@@
for a constant $c>0$ large enough depending on $V$ again by standard interpolation. Using that $\Phi$ is continuous from $L^{2}$ to itself, we get
$\|\Delta v+\mathcal{H}u\|_{L^{2}}\leq\frac{1}{2}\|v\|_{H^{2}}+c\|\Phi\|_{L^{2}\to L^{2}}\|u\|_{L^{2}}$ (2.32)
which completes the proof using
$\|\Delta v\|_{L^{2}}\leq\|\Delta v+\mathcal{H}u\|+\|\mathcal{H}u\|_{L^{2}}$ (2.33)
and the continuity of $\Gamma$ from $L^{2}$ to itself.
@@
Instead of the paracontrolled calculus, one could use the exponential transform as in [17, 21] or the generalized Sturm-Liouville theory as in [14] and references therein to study the operator $\mathcal{H}$. These representation of the operator $\mathcal{H}$ involve a first order term which is problematic to obtain Strichartz inequalities. For example, the exponential transform gives
@@
for $\sigma\in\mathbb{R}$. The previous results give the form domain $\mathcal{D}^{1}=H^{1}$ as for the Laplacian while the domain $\mathcal{D}^{2}=\mathcal{D}_{V}$ depends on $V$.
@@
For $|\sigma|\leq 1+\kappa$, we have $\mathcal{D}^{\sigma}=H^{\sigma}$.
@@
In order to obtain Strichartz inequalities for $\mathcal{H}^{\sharp}$, the comparison with the Laplacian given by the following lemma is needed. This is exaclty where paracontrolled calculus comes into play and improve the naive bound
@@
allows to consider $u(t)=e^{-it\mathcal{H}}u_{0}$ as a solution in $L^{\infty}(\mathbb{R},L^{2})$ which satisfies (1.2) in $\mathcal{D}^{-2}$. In particular, the space $\mathcal{D}^{2}=\mathcal{D}(\mathcal{H})$ does not contain smooth functions thus $\mathcal{D}^{-2}=(\mathcal{D}^{2})^{*}$ is not a space of distributions. To solve the nonlinear equation (1.1) for low regularity initial data, we consider the mild formulation
$u(t)=e^{-it\mathcal{H}}u_{0}-i\lambda\int_{0}^{t}e^{-i(t-s)\mathcal{H}}|u(s)|^{m-2}u(s)\mathrm{d}s$ (3.2)
with $u(0)=u_{0}\in H^{\sigma}$. Since $H^{\sigma}$ is an algebra for $\sigma>\frac{1}{2}$, this gives local well-posedness in $H^{\sigma}$ and global solution for $u\in H^{1}$ with the conservation of energy in the defocusing case $\lambda>0$. In this section, we prove Strichartz inequalities for $\mathcal{H}$ to obtain local well-posedness for low regularity initial data, that is below the algebra condition. In the cubic case without potential, Bourgain obtained local well-posedness in $L^{2}$ thus global solution using the conservation of mass. Even in the cubic case, we have a positive critial threshold $\sigma_{c}>0$ which prevents us from obtaining global solution below $H^{1}$. Our argument is perturbative and relies on Bourgain’s result
@@
$u^{\sharp}(t)=e^{-it\mathcal{H}^{\sharp}}u_{0}-i\lambda\int_{0}^{t}e^{-i(t-s)\mathcal{H}^{\sharp}}\Gamma^{-1}\big{(}|\Gamma u^{\sharp}(s)|^{m-2}\Gamma u^{\sharp}(s)\big{)}\mathrm{d}s$ (3.5)
@@
The loss of derivative here is similar in spirit to the one of Burq, Gérard and Tzvetkov [7] with the Laplacian on an arbitrary manifold and Mouzard and Zachhuber [22] with the Anderson Hamiltonian on compact surfaces. This is due to the fact that our knowledge of the spectral properties of the operator is not precise enough while Bourgain’s method based on Fourier’s series makes full use of the explicit spectrum of the Laplacian.
Proof : We work with the Paley-Littlewood projectors $\Delta_{j}$ on an annulus of size $2^{j}$ in frequencies, see Appendix A. Let $k,j\geq 0$ be fixed. For any $N\geq 1$, we have
@@
using the mild formulation for $e^{i(t-t_{n})\mathcal{H}^{\sharp}}$ with respect to the Laplacian hence